# Parameter-Efficient Subspace Optimization for LLM Fine-Tuning

### v1

@@ G.1 Deterministic Case
G.2 Stochastic Case
We begin by verifying the validity of Assumption 5. As an illustrative case, suppose Opt and UpdateSubspace are implemented by SGD with diminishing step sizes $\{\alpha_{k}\}$ satisfying $\sum_{k}\alpha_{k}<\infty$. Let $\widehat{W}_{k}$ denote the weight after such an update. By the descent lemma (see also [bottou2018optimization, Lemma 4.4]), the expected decrease can be bounded as
For $\alpha_{k}\leq 1/L$, this simplifies to
Taking positive and negative parts, we obtain
Summing over all $k$,
We are now ready to present the proof of our main result, Theorem 5.1.
Proof of Theorem 5.1.
Because $\operatorname*{\mathcal{S}}_{k}$ are subspaces, $P_{\operatorname*{\mathcal{S}}_{k}}$ is a linear operator, which allows the exchangability with $\mathbb{E}$. Therefore, one has
$\begin{split}\mathbb{E}(\|P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k})\|^{2})&=\|\mathbb{E}(P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k}))\|^{2}+\mathbb{V}(P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k}))\\ &=\|(P_{\operatorname*{\mathcal{S}}_{k}}(G_{k}))\|^{2}+\mathbb{E}\left(\|P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k}-G_{k})\|^{2}\right)\\ &\leq\|(P_{\operatorname*{\mathcal{S}}_{k}}(G_{k}))\|^{2}+\mathbb{E}\left(\|\widetilde{G}_{k}-G_{k}\|^{2}\right)\\ &\leq\|(P_{\operatorname*{\mathcal{S}}_{k}}(G_{k}))\|^{2}+C.\end{split}$ (24)
Similar to the deterministic case, We assume the frequency of the full gradient restart is $K$, and consider $k-1\mod K=0$ (i.e., when a full gradient restart occurs). Again, we define
By the property of the full gradient restart, we have
$\begin{split}\mathbb{E}[\ell(\widehat{W}_{k})]&\leq\mathbb{E}[\ell(W_{k-1})]+\left\langle G_{k},\eta_{k}\mathbb{E}[P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k})]\right\rangle+\frac{L}{2}\mathbb{E}[\|\eta_{k}P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k})\|^{2}]\\ &\leq\mathbb{E}[\ell(W_{k-1})]+\eta_{k}\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2}+\frac{L\eta_{k}^{2}}{2}(\|(P_{\operatorname*{\mathcal{S}}_{k}}(G_{k}))\|^{2}+C)\\ &=\mathbb{E}[\ell(W_{k-1})]-\left(\eta_{k}-\frac{L\eta_{k}^{2}}{2}\right)\|P_{\mathcal{S}_{k}}(G_{k})\|^{2}+\frac{L\eta_{k}^{2}}{2}C,\end{split}$ (26)
where the first inequality follows from Assumption 1, and the second follows from the fact that $\operatorname*{\mathcal{S}}_{k}$ is a subspace and (24).
$\mathbb{E}[\ell(W_{i})]\leq\mathbb{E}[\ell(W_{i-1})]+2C_{i},$ (27)
where $2C_{i}$ comes from bounding the scenario where both Opt and UpdateSubspace operate at $i$-th iterate. Then summing up the inequalities (26) and (27) for $i=k+1,\cdots,k+K-1$, and use the fact that $\mathbb{E}[\ell(W_{k})]\leq\mathbb{E}[\ell(\widehat{W}_{k})]+C_{k}$ since $W_{k}$ is obtained by applying Opt to $\widehat{W}_{k}$, we have
By Assumption 1, there exists a constant $C_{\ell}$ so that $C_{\ell}\leq\ell(W)$ for any $W$. Summing up for $k\in\{1,\cdots,T\}$ one has
$\begin{split}\sum_{k=1}^{T}\frac{\eta_{kK+1}}{2}\mathbb{E}[\|P_{\operatorname*{\mathcal{S}}_{kK+1}}(G_{kK+1})\|^{2}]\leq&\frac{LC}{2}\sum_{k=1}^{T}\eta_{kK+1}^{2}+\sum_{k=1}^{T}C_{kK+1}+2\sum_{k=1}^{T}\sum_{i=kK+2}^{(k+1)K}C_{i}\\ &+\mathbb{E}[\ell(W_{K})]-C_{\ell}.\end{split}$ (30)
Finally $\|\mathbb{E}[\ell(W_{K})]-C_{\ell}\|$ is obviously bounded for a fixed $K$, and therefore, one has
By $\sum_{k}^{\infty}\eta_{k}=\infty$ and a contradiction argument, one has $\lim\inf_{k\to\infty}\mathbb{E}[\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|]=0$. Since $\|G_{k}\|\leq dist(G_{k},\operatorname*{\mathcal{S}}_{k})+\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|\leq\delta_{k}+\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|\to\delta$ where $\delta:=\lim_{k\to\infty}\delta_{k}$, the final result follows. ∎
G.3 Additional Theoretical Discussion
There has been a growing body of work establishing exact convergence guarantees for memory-efficient pre-training methods, e.g., [chen2025memory, he2024subspace, robert2024ldadam, zhang2025breaking]. These results typically specify concrete subspace constructions (such as randomized projections), under which the bias term depending on $\delta_{k}$ can be eliminated, provided that the algorithm adopts a carefully tuned hyperparameter schedule, for example a $\beta_{1}$ or learning rate schedule that depends on the iteration horizon. While such schedules are theoretically valid, they often require problem- or time-dependent tuning, which limits their practicality in standard PEFT settings.
To illustrate the compatibility of our framework with such specific cases, we provide a simple case study in which the gradients are deterministic and the subspace $\mathcal{S}_{k}^{\text{FG}}$ is chosen as the top-$r$ SVD subspace of $G_{k}$, as in PESO-LoRA-R. Under this setting, we can establish an exact asymptotic convergence result. We first present the following lemma, which bounds the error introduced by the projected gradients.
@@ G.3 Additional Theoretical Discussion
and define $G_{r}:=U_{r}\Lambda_{r}V_{r}^{\top}$. Then
$\|G-G_{r}\|^{2}\;\leq\;\Bigl(1-\frac{r}{p}\Bigr)\,\|G\|^{2}.$ (33)
@@
$\|G-G_{r}\|^{2}=\sum_{i=r+1}^{p}\sigma_{i}^{2},\qquad\|G\|^{2}=\sum_{i=1}^{p}\sigma_{i}^{2}.$
@@
which implies
$\sum_{i=1}^{r}\sigma_{i}^{2}\;\geq\;\frac{r}{p}\sum_{i=1}^{p}\sigma_{i}^{2}=\frac{r}{p}\,\|G\|^{2}.$
Therefore,
$\|G-G_{r}\|^{2}=\|G\|^{2}-\sum_{i=1}^{r}\sigma_{i}^{2}\;\leq\;\|G\|^{2}-\frac{r}{p}\,\|G\|^{2}=\Bigl(1-\frac{r}{p}\Bigr)\,\|G\|^{2},$
which proves (33). ∎

### v2

@@ G.1 Deterministic Case
G.2 Exact Convergence of Deterministic PESO-LoRA-R
In this section, we present a simple case study in which PESO-LoRA-R admits exact asymptotic convergence. We consider the deterministic setting where the subspace $\mathcal{S}_{k}^{\text{FG}}$ is chosen as the rank-$r$ subspace spanned by the top singular vectors of $G_{k}$, consistent with the construction used in PESO-LoRA-R. Under this setting, we show that the algorithm converges asymptotically to a stationary point without approximation error induced by subspace restriction. This result is consistent with the result in He et al. (2024), which establishes exact convergence for GaLore under deterministic gradients.
We first present the following lemma, which bounds the error introduced by the projected gradients.
@@ G.2 Exact Convergence of Deterministic PESO-LoRA-R
and define $G_{r}:=U_{r}\Lambda_{r}V_{r}^{\top}$. Then the following bounds hold:
$\|G-G_{r}\|_{F}^{2}\;\leq\;\Bigl(1-\frac{r}{p}\Bigr)\,\|G\|_{F}^{2}\quad\text{and}\quad\|G_{r}\|_{F}^{2}\;\geq\;\frac{r}{p}\,\|G\|_{F}^{2}.$ (20)
@@
$\|G-G_{r}\|_{F}^{2}=\sum_{i=r+1}^{p}\sigma_{i}^{2},\qquad\|G\|_{F}^{2}=\sum_{i=1}^{p}\sigma_{i}^{2},\qquad\|G_{r}\|_{F}^{2}=\sum_{i=1}^{r}\sigma_{i}^{2}.$
@@
which implies the second inequality in (20):
$\|G_{r}\|_{F}^{2}=\sum_{i=1}^{r}\sigma_{i}^{2}\;\geq\;\frac{r}{p}\sum_{i=1}^{p}\sigma_{i}^{2}=\frac{r}{p}\,\|G\|_{F}^{2}.$
The first inequality follows immediately from the Pythagorean identity $\|G\|_{F}^{2}=\|G_{r}\|_{F}^{2}+\|G-G_{r}\|_{F}^{2}$:
$\|G-G_{r}\|_{F}^{2}=\|G\|_{F}^{2}-\|G_{r}\|_{F}^{2}\;\leq\;\|G\|_{F}^{2}-\frac{r}{p}\,\|G\|_{F}^{2}=\Bigl(1-\frac{r}{p}\Bigr)\,\|G\|_{F}^{2}.$